OpenStudy (anonymous):

a modulated laser heats a diamond. its variable temperature, in degrees Celsius, is given by f(t)=T sin at. what is the period of the curve?

OpenStudy (mathmale):

A general equation for the sine curve is y=a*sin(b*x). Yours is f(t)=T*sin(a*t). Since the PERIOD of y=a*sin(b*x) is 2Pi/b, the PERIOD of y=T*sin(a*t) is 2Pi/a.

OpenStudy (mathmale):

That's all you need to do. Once more: if y=A*sin (Bx), the period of this function is 2Pi/B.

OpenStudy (anonymous):

but y r u doin 2pi/a ik tht the formula for period is 2pi/b...

OpenStudy (anonymous):

?

OpenStudy (mathmale):

I'm sorry, and I do understand why what I typed could be confusing. Let's start over. If y=A sin Bx, the PERIOD IS 2PI/B. x is your independent variable, and B is the coefficient of that variable. Now, you have f(t) = T sin at. Your independent variable is t. What is the coefficient of t?

OpenStudy (anonymous):

a?

OpenStudy (mathmale):

Yes, and now the PERIOD of YOUR function is \[\frac{ 2\pi }{ coefficient.of.a }\]

OpenStudy (mathmale):

So, what is the period of YOUR function?

OpenStudy (anonymous):

2pi/a